# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v1

@@ 5 Empirical Analysis
Implementing Fourier KArAt. Our implementation of MHSA using the Fourier basis has one hyperparameter, the grid size, $G$, and two configurations, blockwise or universal. Using different permutations of these factors, we can design multiple variants of Fourier-KArAt. We vary our grid sizes from the set ${G=\{1,3,5,7,9,11\}.}$ For each grid value, we perform universal and blockwise updates; see Figure 4. With this formalization, $G_{n}B$ denotes Fourier-KArAt with grid size $n$ and weights updated in blockwise mode, and $G_{n}U$ denotes Fourier-KArAt with grid size $n$ and weights updated in universal mode. See the detailed ablation study of these different variants in the Appendix. We also note that projecting the learned attention rows to a probability simplex degrades the performance; see §B.5. We dispense $\ell_{1}$ projection for the experiments in the main paper. See implementation details in §B.1.
@@
Figure 6: Weight distribution of (top to bottom) ViT-Tiny, ViT-Tiny+Fourier KArAt, ViT-Base, ViT-Base+Fourier KArAt. The columns (left to right) represent weights at initialization, at epoch 50, at epoch 100, and their superposition.
Figure 7: Spectral decomposition of the attention matrix for ViT-Tiny on CIFAR-10 dataset with vanilla softmax attention and our learnable Fourier attention.
Figure 8: Loss landscape for ViT-Tiny and ViT-Base (the smallest and the largest model) along the two largest principal component directions of the successive change of model parameters. The top row provides a 3D-visulaization of the loss surface including the local optima and saddle points. The bottom row shows the loss contours along with the trajectory of the optimizers.
5.1 Training Performance and Model Quality
@@
Table 1 presents the performance of the best-performing Fourier KArAt variants. We also present the training loss and test accuracy curves on all three datasets; see Figure 5. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT Tiny+Fourier KArAt), the loss curve flattens faster. From Figure 5 we infer that for the majority of the scenarios involving Fourier-KArAt. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process difficult to manage, which we try to investigate in the rest of the paper using several analytical studies. To conclude, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K it achieves almost similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models and the accuracy differences increase.
Comparing the Top-1 accuracies from Table 1 across blockwise and universal modes on the datasets does not have enough evidence to support the claim that one mode is better than the other. These tabulated accuracies are the results of applying the two variants on only two grid sizes.
Figure 9: Visualizing attention maps for Vit-Tiny. The original images used for inference are on the left, and on the right, we show the attention score map and the image regions of the dominant head (Top row: Fourier KArAt, bottom row: Vanilla MHSA).
5.2 Analysis of Performance
In this Section, we dissect the performance of Fourier KArAt through a diverse set of analyses.
@@ 5.2 Analysis of Performance
(i) Distribution of the Weights. (Zhang et al., 2022) considered an invariant measure perspective to describe the training loss stabilization of the neural networks. Based on this idea, we study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 6 shows the distribution of the weights of these models during different training phases. The evolution of the weights’ distributions for respective models remains invariant, and we can guarantee the convergence of their loss values; also, see Figure 5. However, as explained by (Zhang et al., 2022), this perspective cannot shed light on the generalization capacity and structural differences between different neural networks.
(ii) Spectral Analysis of Attention. Although by studying the distribution of the weights we cannot comment on the generalizability, from Table 1 we realize all the KArAt variants have more parameters than their vanilla counterparts. Therefore, it would be interesting to see how their attention matrices behave. As discussed in §4, the attention matrices in vanilla MHSA have a low-rank structure. Following that study, we investigated if Fourier KArAt’s learned attention matrices are also low-rank. We use all 3 heads in the last encoder block of ViT-Tiny on 5 randomly sampled images from the CIFAR-10 validation set. There are a total of 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order. Let $\sigma_{i}$ be the $i^{\rm th}$ singular value across all heads and samples (it is permutation invariant). For each $i\in[197],$ we plot $[\ln(\sigma_{i}^{\min}),\ln(\bar{\sigma_{i}}),\ln(\sigma_{i}^{\max})]$, where $\bar{\sigma_{i}}$ is the average of $i^{\rm th}$-indexed singular value across all samples and heads. From Figure 7, we observe that vanilla softmax attention has significantly larger singular values than its KArAt variant, $G_{3}B$. Fourier KArAt with more parameters shows a faster drop in significant singular values. This observation indicates that Fourier KArAt variants are highly overparameterized and have a much lower rank structure compared to the vanilla ViTs.
(iii) Loss Landscape Analysis. To understand the reason behind the slower convergence in the later training phase and the lower generalizability of KArAts, we visualize the loss landscape of the networks. We plot the loss surface along the directions in which the gradients converge. Following (Li et al., 2018), we perform principal component analysis (PCA) of the change in parameters over training progression to understand the major directions of parameter convergence. Considering the fully trained model as the minimum in the loss hyperplane and the two principal component directions as $X$ and $Y$ axes, we plot the loss values over the validation set of CIFAR-10 along the $Z$ axis for ViT-Tiny and ViT-Base for the vanilla MHSA and KArAt ($G_{3}B$ and $G_{1}B$, respectively). These visualizations in Figure 8 show that having Fourier KArAt in ViT architectures significantly impacts the smoothness of the loss surfaces.
@@
ViT-Tiny with the least parameters has the smoothest loss landscape and modest generalizability. In contrast, ViT-Tiny+Fourier KArAt’s loss landscape is spiky; it indicates the model is full of small-volume minima (Huang et al., 2020). However, the model is modest in the number of parameters, so the gradient descent optimizer can still find an optimized model with better generalizability than the vanilla ViT-Tiny; hence it gains a better test accuracy; see Table 1. ViT-Base, however, has more parameters than Tiny and its loss surface is much spikier than ViT-Tiny. Finally, the loss surface of ViT-Base+Fourier KArAt is most spikey, making it a narrow margin model with sharp minima in which small perturbations in the parameter space lead to high misclassification due to their exponentially larger volume in high dimensional spaces; see Figure 8(d). Moreover, ViT-Base+$G_{3}B$ has 14 times more parameters than ViT-Tiny+$G_{1}B$. With learnable activations, gradient descent optimizers fail to find the best-optimized model, as small differences in the local minima translate to exponentially large disparities. The slow convergence in the later phases of the models is bacause the increasing number of local minima can slow down convergence.
@@
(iv) Attention Visualization. MHSA in ViTs captures region-to-region interaction in images. DINO (Caron et al., 2021) provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is not directly suitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with vanilla MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figure 9. Fourier KArAts tend to concentrate high interaction scores on the concerned object’s dominant features compared to having interactions throughout the object regions like traditional ViTs (Dosovitskiy et al., 2020; Touvron et al., 2021).
6 Conclusion and Future Direction
This paper proposes a learnable MHSA for vision Transformers that can theoretically operate on any choice of basis. The central design theme of this learnable attention module is based on the famous Kolmogorov-Arnold representation theorem (Kolmogorov, 1956) and the Kolmogorov-Arnold network proposed by (Liu et al., 2025).
This endeavor raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalization capacity. In the future, with the advent of large multimodal models (LMMs), we plan to extend this study beyond ViTs for image classification and check the resilience of KArAt in the language processing domain and on open-source LMMs.
KANs are parameter efficient—this claim does not hold for learnable KArAt selft-attention design. However, our spectral analysis and loss-landscape experiments suggest that overparameterized KArAts’ weights lie in a much lower-dimensional subspace. Nevertheless, how to navigate through KArAts’ parameter-heavy loss landscape to find the optimized model with better generalizability remains an open question that we encourage the community to focus on. Finally, we note that the data specificity of MHSA plays a key role in their high performance. Therefore, we aim to explore its scope with KArAts with the possibility of pitfalls due to long-range dependency (Park & Kim, 2022).
Appendix A Solution to Problem (3)
@@
The stationarity and complementary slackness conditions give $x_{i}-y_{i}+\lambda\geq 0,\ (x_{i}-y_{i}+\lambda)x_{i}=0$. If $\lambda\geq y_{i}$, then $x_{i}=0$. Otherwise, we have that $\sum\limits_{i=1}^{m}x_{i}=\sum\limits_{i=1}^{m}(y_{i}-\lambda)=1$ implying $\lambda=\frac{1}{m}(\sum\limits_{i=1}^{m}y_{i}-1)$.
@@
Note that, in Algorithm 1 one can replace $y$ with $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]$ so that the constraint ${\|\widehat{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\widehat{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ is satisfied for the low-rank attention matrices in Fourier-KArAt. This algorithm is also kept as an optional sub-process for Algorithm 2.
Appendix B Addendum to Empirical Analysis

### v2

@@ 5 Empirical Study
Implementing KArAt. We provide the implementation details in §B.1; see §B.1.1 for baselines and datasets. We integrate 5 wavelet bases, rational basis, Fourier basis, and 3 different base activation functions in our learnable MHSA design; see §B.1.3. Fourier KArAt performs the best, and we report it in the main paper. It has 1 hyperparameter, the grid size, $G$, and 2 configurations, blockwise or universal. We vary grid sizes from the set ${G=\{1,3,5,7,9,11\}.}$ For each value, we perform universal and blockwise updates; see Figure 2. With this formalization, $G_{n}B$ and $G_{n}U$ denote Fourier-KArAt with grid size $n$ and weights updated in blockwise and universal mode, respectively. See the ablation study of these variants in §B.4. We dispense the $\ell_{1}$ projection, as projecting the learned attention vectors to a probability simplex degrades the performance; see §B.3.5.
Table 3: Performance comparison of Fourier-KArAt and softmax attention on various ViT architectures.
Model CIFAR-10 ImageNet-1K
Acc.@1 Acc.@5 Acc.@1 Acc.@5
ConViT-Tiny 71.36 97.86 57.91 81.79
+ $G_{3}B$ 75.57 98.61 56.57 80.75
+ $G_{3}U$ 74.51 98.63 56.51 80.93
Swin-Tiny 84.83 99.43 76.14 92.81
+ $G_{3}B$ 79.34 98.81 73.19 90.97
Table 4: Fourier KArAt on object detection and instance segmentation tasks on MS COCO [48] dataset. The header Box and Mask refer to detection and segmentation tasks, respectively.
Model (Initialization) Box Mask
AP AP50 AP75 AP AP50 AP75
ViT-Det-Base (ImageNet-1K) 32.34 49.16 34.43 28.35 46.04 29.48
+ $G_{1}B$ 22.48 36.82 23.13 20.11 34.09 20.46
+ $G_{1}U$ 26.68 43.25 27.88 24.01 40.21 24.66
ViT-Det-Base (Random) 15.28 26.88 15.22 13.39 24.34 12.86
+ $G_{1}B$ 10.32 18.77 10.05 8.94 16.51 8.50
+ $G_{1}U$ 10.99 20.04 10.82 9.69 17.99 9.25
(a) ViT-Tiny
(b) ViT-Tiny+Fourier KArAt
(c) ViT-Base
(d) ViT-Base+Fourier KArAt
@@
Figure 3: Weight distribution of ViT-Tiny and -Base with traditional MHSA and Fourier KArAt. The columns (left to right) represent weights at initialization, epoch 50, and epoch 100.
5.1 Training Performance, Model Quality, Transferability, and Generalizability
@@
(i) Image Classification. Table 2 presents the performance of the best-performing Fourier KArAt variants; see Figure 7 for the training loss and test accuracy curves. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT-Tiny+Fourier KArAt), the loss curve flattens faster; we infer that for most of the Fourier-KArAt scenarios. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process challenging to manage, which we try to investigate in the rest of the paper using several analytical studies. Overall, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K, it achieves a similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models, and the accuracy differences increase as we tackle larger base models.
(ii) Performance on other ViTs for Image Classification. We fused KArAt with modern ViT-based architectures, ConViT [20] and Swin Transformer [49]; §B.1.4 provides their implementation details. Table 4 shows that Fourier KArAt outperforms ConViT on CIFAR-10, experiences a modest drop on ImageNet-1K, and shows lower performance for hierarchical models, such as Swin, which requires a careful investigation.
(iii) Transferability: Image Classification. Unlike ViTs, which were pre-trained on large datasets and then transferred to various mid-sized or small image recognition benchmarks [24], which helped them achieve state-of-the-art results, KArAts were never pre-trained with such datasets to showcase their full potential. Due to limited computing resources, we could not perform large-scale training. We investigate the KArAts’ transfer capability by fine-tuning on smaller datasets like SVHN [54], Oxford Flowers 102 [59], and STL-10 [18] from their ImageNet-1K pre-trained weights; see dataset details in §B.1.1. Table 2 shows KArAts transfer well across all datasets and have comparable performance with vanilla ViTs, even when their ImageNet-1K performance was not equivalent.
(iv) Generalizability: Detection & Segmentation. To understand the robustness of the Fourier KArAt, we employ ViT backbones for detection and segmentation. We choose Mask RCNN [33]-based framework for this purpose and employ a feature pyramid network [47] based on ViT-Det [46]; it is non-trivial to implement Fourier KArAt, see §B.1.5. We trained ViT-Det [46] using ViT-Base backbone on the MS COCO [48] dataset for 50 epochs with two settings: (i) fine-tuning on the ImageNet-1K pre-trained weights on traditional softmax attention, and (ii) random initialization. Table 4 shows that for the fine-tuning on ImageNet pre-trained weights, Fourier-KArAt shows $\sim 4-13\%$ gap in average precision (AP) from its vanilla variant; their qualitative performance in Figure 9 is similar. Overall, for ViT-Det, Fourier-KArAt performs inferiorly to its conventional counterpart in detection and segmentation. Interestingly, Figure 10 shows, vanilla ViT-Det using softmax activation reaches its peak performance quickly (within 12 epochs), Fourier-KArAt delays in achieving it (within 30 and 45 epochs for $G_{1}B$ and $G_{1}U$, respectively), proving the incompatibility of softmax-based weights initialization. It is, however, still a good initialization for KArAt as the performance is better than random initialization.
Key Takeaways. (a) The Top-1 accuracies from Tables 2 and 2 show Blockwise configuration is more desirable over Universal for image classification. (b) From Table 2, we can infer that the Fourier KArAt transfers well in fine-tuning tasks. However, the transferability of hyperparameters, e.g., grid size $G$, across datasets remains an open question. (c) KArAt generalizes well. For random initialization in detection, the vanilla ViT-Det has an advantage over KArAt, and thus, we see a small performance gap. The best KArAt hyperparameters for this task are yet to be found. Overall, KArAt shows significant potential if the incompatibilities are properly addressed.
Figure 4: 3D-visualization of Loss landscape for ViT-Tiny and ViT-Base along the two largest principal component directions of the successive change of model parameters. KArAt’s loss landscapes are significantly less smooth than traditional attention; spiky loss landscapes are undesirable in terms of optimization stability and generalizability of the resulting model. See Figure 11 for the loss contours and the optimizer trajectory.
5.2 Performance Analysis
We dissect the performance of Fourier KArAt using the following tools:
@@ 5.2 Performance Analysis
(i) Distribution of the Weights. [84] considered an invariant measure perspective to describe the training loss stabilization of the neural networks. We adopt this idea to study the distribution of the weights of the smallest and largest models of the ViTs, Tiny and Base, and their Fourier KArAt variants. Figure 3 shows the distribution of the weights of these models during different training phases—The evolution of the weights’ distributions for respective models remains invariant. Based on this observation, from [84], we can guarantee the convergence of loss values of all models; also, see Figure 7. However, as mentioned in [84], this perspective cannot comment on the generalization capacity and structural differences between different neural networks.
@@
(ii) Loss Landscape. We visualize the loss landscape of KArAts to understand why they converge slowly in the later training phase, and why their generalizability does not scale with larger models. Following [45], we plot the loss surface along the directions in which the gradients converge; see §B.1.6 for implementation. These visualizations in Figure 4 show that Fourier KArAt in ViT architectures significantly impacts the smoothness of the loss surfaces. ViT-Tiny, with the fewest parameters, has the smoothest loss landscape and modest generalizability. In contrast, ViT-Tiny+Fourier KArAt’s loss landscape is spiky; it indicates the model is full of small-volume minima [35]. However, the model is modest in the number of parameters, so the gradient descent optimizer can still find an optimized model with better generalizability than the vanilla ViT-Tiny; hence, it gains a better test accuracy; see Table 2. ViT-Base, however, has more parameters than Tiny, and its loss surface is much spikier than ViT-Tiny. Finally, the loss surface of ViT-Base+Fourier KArAt is most spiky, making it a narrow margin model with sharp minima in which small perturbations in the parameter space lead to high misclassification due to their exponentially larger volume in high dimensional spaces; see Figure 4(d). Moreover, ViT-Base+$G_{3}B$ has 14 times more parameters than ViT-Tiny+$G_{1}B$. With learnable activations, gradient descent optimizers fail to find the best-optimized model, as small differences in the local minima translate to exponentially large disparities. The increasing number of local minima in the later phases of the training causes the slow convergence.
@@
(iii) Attention Visualization. MHSA in vanilla ViTs captures region-to-region interaction in images. DINO [15] provides an innovative way to explain the dominant regions in an image that contribute towards the inference decision. It maps the self-attention values of the CLS token at the last encoder layer, which is directly responsible for capturing the information related to the class label. While the computer vision community predominantly uses this attention visualization, it is unsuitable for our case. Unlike softmax, Fourier KArAt does not inherently ensure learned attention values in $[0,1]$, as we dispense the $\ell_{1}$ projection. Therefore, we cannot interpret its performance similarly. Nonetheless, pre-softmax or pre-KArAt values in the attention matrix $\mathcal{A}^{i,j}$ are supposed to capture token-to-token interactions, and we adapt the attention maps to ignore the negative values for the sake of visualization. We use the ImageNet-1K trained models of ViT-Tiny (with traditional MHSA and Fourier KArAt) and plot the attention maps of the last layers in Figures 5 and 12.
Figure 5: Vit-Tiny Attention map visualization. Original images for inference (the left), the attention score (middle), and image regions of the dominant head (Top row: Fourier KArAt, bottom row: traditional MHSA).
6 Discussion, Limitations, and Conclusion
The learnable MHSA design raises many interesting questions regarding KArAt’s computing scalability and the resulting models’ generalizability. Below, we discuss their potential positive ($+$) and negative (--) implications and encourage researchers to elaborate on them.
Designing KArAt self-attention is not parameter efficient and they are memory hungry (--). Having learnable activations in the MHSA causes a memory explosion on the present GPUs. Hence, calculations with the full-rank operator, ${\Phi\in\mathbb{R}^{N\times N}},$ are prohibitively expensive, regardless of the basis. Present GPUs and CUDA functions are not optimized for KAN implementation as noted by many before us; see [80, 66, 65, 78, 61]. Due to this, although parameter increase is not always significant (Table 2), optimized Fourier KArAt variants utilize 2.53×2.5-3$\times$2.5 - 3 × more GPU memory compared to traditional softmax attention. We propose a low-rank approximation to reduce memory requirements to a feasible level (Table 12), but researchers should explore alternative techniques. Interestingly, while there is a significant training time discrepancy between vanilla ViTs and Fourier-KArAts, their inference speeds are comparable on the present compute interface and hardware configuration; see Figure 8 in §B.2. Hence, we hope the future holds more efficient system-level optimization recipes for training KArAt. Additionally, we witnessed KArAts’ inconsistent training behavior with parameter changes in §B.3.4. Each ViT model with particular Fourier-KArAt variants has a typical $G$ value that brings out its best performance; there is no universal $G$ to follow.
KArAts show a better attention score ($+$) and transfer learning capability ($+$) for all ViTs, but their generalizability does not scale with larger ViTs (--). Fourier KArAt concentrates high interaction scores on the object’s dominant features compared to having interactions throughout the object regions like vanilla ViTs [24, 73]. They exhibit decent transfer learning and generalizability. The distribution of weights guarantees that their training losses stabilize. Hence, by [84], we can postulate that first-order optimization algorithms (e.g., ADAM) can optimize KArAt’s loss landscape. But KArAt’s generalizability does not scale with larger models due to high parameterization, and [84] cannot explain this. We encourage researchers to investigate these models’ local or global optimality and generalizability through rigorous theoretical and empirical studies.
Overparameterized KArAts’ weights lie in a much lower-dimensional subspace ($+$). Spiky loss landscape with local minima is probably the major impediment towards KArAt’s scalability; see Figure 4. Modern transformers are highly over-parameterized, and their weight matrices show low-rank structures [41]; spectral analysis (§B.5) of KArAt shows its attention matrices possess a better low-rankness than traditional attention (Figure 13). Considering this, properly guided parameter reduction in KArAt can result in an optimized model that generalizes well at scale.
KArAts can provide better interpretability for smaller models ($+$). KANs are claimed to be more interpretable and accurate than MLPs due to their flexible choice of univariate functions [51]; it is also evident from KArAts’ attention maps. ViT-Tiny+KArAts outperforms vanilla ViT-Tiny on small-scale datasets, and performs at par on ImageNet-1K (Table 2) and transfer learning (Table 2). In the future, KArAt’s learnable activation makes a case for investigating it to add more explainability in vision tasks and can find more interpretability in smaller models trained in a limited data scenario.
Conclusion.  Across our extensive benchmarking, KArAt outperforms traditional MHSA in smaller ViTs; results on larger models are mixed. Additionally, we note that the computing requirement impedes KArAt’s performance—training and fine-tuning these models are expensive. However, learnable activations can make a substantial difference in a model’s interpretability, and these limitations are not KArAt’s weaknesses. At its early stage, KArAt’s decent performance on diverse tasks shows its potential. In the future, one can extend this study beyond ViTs and check the resilience of KArAt in language processing and on large multimodal models [2, 79, 72, 13, 11, 3, 71, 4].
Appendix
Organization. We organized the Appendix as follows: §A contains the solution to problem (4). We also provide the pseudocode to implement Fourier-KArAt in ViT’s encoder block; see Algorithm 2. §B discusses additional numerical results, including implementation details, hyperparameter tuning, computation time analysis, diverse ablation studies, and different variants of Fourier KArAt. Finally, we concluded §B with the spectral analysis of the attention matrices for traditional and learnable attention.
Appendix A A Sparse Approximation Problem and its Solution
Let $y\in\mathbb{R}^{m}$ be a given vector. If we want to approximate $y$ with a vector, $x\in\mathbb{R}^{m}$, with positive components and $\|x\|_{1}=1$, we can write the constrained optimization problem as:
$\begin{cases}&x^{\star}=\arg\min_{x\in\mathbb{R}^{m}}\frac{1}{2}\|x-y\|^{2}\\ &\text{subject to\;\;}x_{i}\geq 0\;\text{for}\;i\in[m],\text{and}\;\|x\|_{1}=1\text{.}\\ \end{cases}$ (4)
@@ Appendix A A Sparse Approximation Problem and its Solution
The stationarity and complementary slackness conditions give, $x_{i}-y_{i}+\lambda\geq 0,\ (x_{i}-y_{i}+\lambda)x_{i}=0$. If $\lambda\geq y_{i}$, then $x_{i}=0$. Otherwise, we have that $\sum\limits_{i=1}^{m}x_{i}=\sum\limits_{i=1}^{m}(y_{i}-\lambda)=1$ implying $\lambda=\frac{1}{m}(\sum\limits_{i=1}^{m}y_{i}-1)$.
Algorithm 1 Projection on $\ell_{1}$ ball [19]
  Input: $y\in\mathbb{R}^{m}$
  Sort: $y$ as $y_{(1)}\geq y_{(2)}\geq...\geq y_{(m)}$
  Calculate: $\rho\triangleq\max\limits\Big{\{}i\in[m]|\ y_{(i)}-\frac{1}{i}\bigg{(}\sum\limits_{j=1}^{i}y_{(j)}-1\bigg{)}\Big{\}}$
  Set: $\lambda=\frac{1}{\rho}\bigg{(}\sum\limits_{i=1}^{\rho}y_{(i)}-1\bigg{)}$
  Output: $x^{\star}=(y-\lambda)_{+}\triangleq\max{(y-\lambda,0)}$
@@
Note that, in Algorithm 1, one can replace $y$ with $\widehat{\Phi}^{i,j}[({\cal A}^{i,j})_{k,:}^{\top}]$ so that the constraint ${\|\widehat{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\widehat{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ is satisfied for the low-rank attention matrices in Fourier-KArAt. This algorithm is also kept as an optional sub-process for Algorithm 2.
Appendix B Addendum to Empirical Study